OpenStudy (anonymous):

Circle P is tangent to the x-axis and the y-axis. If the coordinates of the center are (r, r), find the slope of the line through the origin and the center.

OpenStudy (raden):

the slope which passes through the points (r,r) and (0,0) is m = (y2 - y1)/(x2 - x1) = (r - 0)/(r - 0) = r/r = 1